Suppose a country's real GDP is $14 trillion and that population is 200 million. Instructions:

Enter your answers as whole numbers.

a. What is this country's real GDP per capita? $ 70000 Suppose that during the next 10 years, real GDP triples and population doubles in the country.

b. After 10 years have passed, what will be this country's real GDP per capita?
Answer #1

ANSWER:

A) Real gdp per capita = Real gdp / population = 14,000,000,000,000 / 200,000,000 = 70,000

so the real gdp per capita is $70,000

B) Real gdp 10 years later = 14,000,000,000,000 * 3 = 42,000,000,000,000

population after 10 years = 200,000,000 * 2 = 400,000,000

Real gdp per capita = Real gdp 10 years later / population after 10 years = 42,000,000,000,000 / 400,000,000 = 105,000

so the real gdp per capita is $105,000